Trevor launched a rocket in the soccer field. it landed 22 feet above the ground in a tree. if the rocket and trevor are 25 feet apart, what is the distance between trevor and the tree?
Answer by ewatrrr(24785)


Hi
Trevor launched a rocket in the soccer field, landing 22 feet above the ground in a tree
if the rocket and trevor are 25 feet apart, find distance between trevor and the tree
Let x represent distance between trevor and the tree
Using the Pythagorean Theorem
x^2 = 25^2 - 22^2
x = sqrt(141)= 11.8743 ft